# Availability-aware Sensor Fusion via Unified Canonical Space for 4D Radar, LiDAR, and Camera

### v1

@@ 5 Conclusion
This paper introduces availability-aware sensor fusion (ASF), which addresses sensor availability challenges in AD by transforming features into a unified canonical space through UCP and CASAP-PN. Our approach maintains computational efficiency ($O(N_{s}^{2}N_{p})$) while providing robust fusion for sensor degradation or failure. The proposed sensor combination loss further enhances robustness by optimizing across all possible sensor combinations. Experiments on the K-Radar dataset demonstrate significant improvements over SOTA methods (9.7% in $AP_{BEV}$ and 20.1% in $AP_{3D}$ at IoU=0.5), with consistent performance across various weather conditions and sensor combinations.
@@
Appendix. The appendix is organized as follows. Section A provides a detailed explanation of the sensor fusion framework, including specific implementations of each component. Section B presents additional t-SNE visualizations that further demonstrate the effectiveness of the proposed method. Section C includes additional performance results across various weather conditions and sensor combinations.
@@ Appendix A Details of Sensor Fusion Framework
Figure 5: Visualization of feature representations through t-SNE [1] at different stages of the proposed ASF for the ‘Bus or Truck’ class. ‘UCP’ and ‘CASAP-PN’ denote unified canonical projection and cross-attention across sensors along patches with post-feature normalization, which are the two main components of the proposed ASF.
As shown in Fig. 4, the sensor fusion framework is divided into three stages: (1) sensor-specific encoders [45, 35, 39, 46, 12, 15] that extract feature maps (FMs) from different types of multi-modal sensors (e.g., camera, LiDAR, and 4D Radar), (2) a fusion method that integrates these FMs, and (3) a task-dependent detection head. Stage (1) has evolved independently for each sensor [47, 48, 12], while stage (3) employs a task-dependent method (e.g., YOLO head [49] for object detection) regardless of sensor type. Therefore, in this work, we propose a fusion method focusing on stage (2) rather than stages (1) and (3). Details of both existing and proposed fusion methods are illustrated in Fig. 2. In this section, we describe both the sensor-specific encoders and the detection head of the proposed availability-aware sensor fusion (ASF).
@@ A.1 Sensor-specific Encoders
Sensor-specific encoders extract feature maps (FMs) from each sensor’s data. Because different sensors provide data in different modalities—cameras produce 2D RGB images, LiDAR generates 3D point clouds, and 4D Radar outputs low-resolution tensors with power values—research on these encoders has evolved independently for each sensor type.
@@
For ASF, we adopt established backbone architectures for each sensor and pre-train these backbones before ASF training. Each encoder is trained for 11 epochs following the K-Radar [12] training protocol. During ASF training, we freeze all sensor-specific encoders to ensure that only the ASF components (UCP and CASAP-PN) are learned. Details for each encoder are as follows:
@@
Camera. Camera-based sensor encoders for autonomous driving typically require transforming front-view (FV) images into a bird’s-eye-view (BEV) format for 3D object detection. Current approaches can be broadly categorized into two methods: (1) learnable BEV transformations such as Lift-Splat-Shoot (LSS) [35, 18, 17], or (2) maintaining FV while using transformer decoders [24] that incorporate depth information as positional embeddings in BEV space. Although methods like PETR [24] have improved 3D object detection by injecting depth embeddings, they inevitably increase computational complexity due to the additional depth dimension. These two approaches form the foundation of camera-based fusion [20, 19, 18, 17, 5], commonly seen in deeply coupled fusion (DCF) and sensor-wise cross-attention fusion (SCF).
@@
To develop a computationally efficient method, we opt for the LSS approach and adopt BEVDepth [38], which supervises the LSS depth estimation using LiDAR point clouds. For encoding images, we use the Swin-S model of SwinTransformer as the 2D backbone, achieving 83% top-1 accuracy on ImageNet [50]. We resize K-Radar’s front camera images to $704\!\times\!256$ pixels for encoding. The encoded 2D feature maps are projected into BEV feature maps via LSS with a grid size of 0.2 m. To match the 0.4 m grid size used in K-Radar, we add a convolutional layer with stride 2. During training, we employ the depth information projected from LiDAR point clouds onto the image and, inspired by [51], we use the LiDAR backbone’s (SECOND) BEV feature maps for distillation. As shown in Fig. 3, camera images have limited utility in adverse weather, so we train the camera encoder only on normal or overcast scenes (sequences 1–20) from K-Radar.
@@
A limitation of this approach is that promising stereo or multi-view detection models cannot be fully leveraged, because the K-Radar dataset [12]—the only one offering camera, LiDAR, and 4D Radar data in adverse weather—provides object annotations only for the front view. Moreover, unlike the KITTI dataset with its longer stereo baseline of 0.54 m, K-Radar uses waterproof ZED cameras with a baseline of just 0.12 m, limiting stereo-based detection for distant objects. Consequently, we employ a monocular camera method despite inherent performance constraints.
@@
LiDAR. LiDAR provides precise depth information in the form of dense point clouds, often used as a reference for labeling other sensors [12, 52]. We adopt the SECOND [39] model as the LiDAR encoder, which is open-source in K-Radar and widely used in various benchmarks. Leveraging LiDAR’s high spatial resolution, we set the voxel size to 0.05 m and apply multiple layers of sparse convolution to generate BEV feature maps with a 0.4 m grid size.
@@
4D Radar. Unlike LiDAR, which uses Time-of-Flight (ToF) measurement, Radar detects spatial information through signal processing, which can introduce sidelobe noise. Traditional 3D Radar typically filters out noisy measurements to produce a sparse point cloud focusing on object presence [11]. In contrast, 4D Radar captures additional elevation information, yielding higher-density point clouds [53]. This property is exemplified by K-Radar [12, 14]. We adopt RTNH [12], the baseline provided in K-Radar, as our 4D Radar encoder, following the original network settings and design.
@@ A.2 Detection Head
Following the K-Radar baseline, we use an SSD detection head [40, 41] for the 3D object detection task. Our implementation uses two anchor boxes per object class, oriented at $0^{\circ}$ and $90^{\circ}$. Each anchor box is parametrized by eight values: the 3D center coordinates ($x_{c}$, $y_{c}$, $z_{c}$), 3D size ($x_{l}$, $y_{l}$, $z_{l}$), and orientation ($\cos(\text{yaw})$, $\sin(\text{yaw})$), as in [41].
To address the severe class imbalance in object detection (negative samples far outnumber positives), we use the focal loss [54] for classification. For bounding box regression, we use smooth L1 loss to measure differences between predictions and ground truth, improving training stability. During inference, we select the class with the highest logit score for each proposal, then apply non-maximum suppression to remove redundant overlapping detections, producing the final predictions.
Figure 6: Qualitative results of ASF with C+L+R (C: Camera, L*: damaged LiDAR, R: 4D Radar). We show results under (a–b) normal and (c–d) adverse weather conditions. Each grid shows the camera image, LiDAR point cloud, 4D Radar tensor, and a sensor attention map (SAM) illustrating attention scores from CASAP-PN (except for (a) and (c), which only show sensor data for ground truth visualization). In the SAM, red, green, and blue represent attention to Camera (C), LiDAR (L), and 4D Radar (R), respectively. For example, a predominantly blue SAM indicates that 4D Radar has the highest attention in that scene. The bottom-left corner of each grid shows attention score proportions (C/L/R[%]). Yellow and red boxes denote predictions and ground truth, respectively.
@@ B.1 Additional t-SNE Visualizations
This section provides additional feature visualizations to complement those in Fig. 1, which only shows results for the ‘Sedan’ class. Fig. 5 presents t-SNE [1] visualizations of the ‘Bus or Truck’ class at different stages of the ASF. These visualizations further demonstrate how ASF unifies feature representations from different sensors.
Consistent with Fig. 1, we observe three main trends in the t-SNE visualizations. First, the initial features from the sensor-specific encoders are scattered without clear alignment across different sensors. Next, after the unified canonical projection (UCP), sensor-specific features become better aligned with the fused features. Finally, following CASAP-PN, features from available sensors form cohesive clusters (in this case, LiDAR and 4D Radar).
These findings underscore the effectiveness of ASF in creating a unified canonical space for multi-sensor features. Note that in K-Radar, the ‘Sedan’ class has 37,213 instances, whereas the ‘Bus or Truck’ class has 7,449 instances, leading to somewhat sparser data in Fig. 5 compared to Fig. 1.
@@ B.2 Details of t-SNE Visualizations
To visualize object-specific features in each sensor’s feature map (FM), we use RoIAlign [55] to crop features within bounding boxes. RoIAlign extracts fixed-size feature maps, making it suitable for t-SNE (which requires consistent dimensions). We then visualize these cropped features in the unified canonical space. For those extracted before projection (where channel dimensions may differ), we downsample as needed to ensure consistent feature dimensions.
Appendix C Additional Detection Results
This section presents additional detection results for ASF. First, we provide qualitative examples where the LiDAR sensor is damaged. Then, we show quantitative performance, including $AP_{3D}$ and $AP_{BEV}$ for both ‘Sedan’ and ‘Bus or Truck’ at IoU=0.3 and 0.5 on K-Radar [12] benchmark v2.0. These complement Tab. 3, which only reports $AP_{3D}$ at IoU=0.3.
C.1 Additional Qualitative Results on Damaged LiDAR
@@
When LiDAR surfaces are damaged or contaminated, LiDAR measurements can be lost [56]. In adverse conditions such as heavy snow, accumulation on the LiDAR sensor can block forward measurements, as illustrated by the red dashed line in Fig. 6-(d), even though vehicles are present. We also simulate total loss of forward LiDAR returns from surface damage (e.g., broken sensor), as seen in Fig. 6-(b). Despite missing LiDAR data, ASF still detects all objects using only camera and 4D Radar inputs. As reported in Tab. 3, ASF with C+L+R often performs similarly or even better than C+R alone. The sensor attention map (SAM) in Fig. 6-(b) confirms the increased reliance on camera or 4D Radar when LiDAR is unavailable.

### v2

@@ 5 Conclusion
This paper introduces availability-aware sensor fusion (ASF), which addresses sensor availability challenges in autonomous driving by transforming features into a unified canonical space through UCP and CASAP. Our approach maintains computational efficiency ($O(N_{q}N_{s})$) while providing robust fusion for sensor degradation or failure. The proposed sensor combination loss further enhances robustness by optimizing across all possible sensor combinations. Experiments on the K-Radar dataset demonstrate significant improvements over SOTA methods (9.7% in $AP_{BEV}$ and 20.1% in $AP_{3D}$ at IoU=0.5), with consistent performance across various weather conditions and sensor combinations.
@@
Appendix
The appendix is organized as follows. Section A provides detailed explanations of the sensor fusion framework, including specific implementations of sensor-specific encoders and the detection head. Section B presents additional t-SNE visualizations that demonstrate the effectiveness of the unified canonical space. Section C analyzes attention score distributions across different weather conditions and distance ranges. Section D includes comprehensive performance results across various sensor combinations and alternative stereo camera configurations.
@@ Appendix A Details of Sensor Fusion Framework
As shown in Fig. 4, the sensor fusion framework is divided into three stages: (1) sensor-specific encoders (swin, lss, second, pvrcnn, kradar, rtnhp) that extract feature maps (FMs) from different types of multi-modal sensors (e.g., camera, LiDAR, and 4D Radar), (2) a fusion method that integrates these FMs, and (3) a task-dependent detection head. Stage (1) has evolved independently for each sensor (vit, pointpillars, kradar), while stage (3) employs a task-dependent method (e.g., YOLO head (yolo) for object detection) regardless of sensor type. Therefore, in this work, we propose a fusion method focusing on stage (2) rather than stages (1) and (3). Details of both existing and proposed fusion methods are illustrated in Fig. 1. In this section, we describe both the sensor-specific encoders and the detection head of the proposed availability-aware sensor fusion (ASF).
@@ A.1 Sensor-specific Encoders
Sensor-specific encoders extract FMs from data of each sensor. Because different sensors provide data in different modalities—cameras produce 2D RGB images, LiDAR generates 3D point clouds, and 4D Radar outputs low-resolution tensors with power values—research on these encoders has evolved independently for each sensor type.
@@
For ASF, we adopt established backbone architectures for each sensor and pre-train these backbones before ASF training. Each encoder is trained for 11 epochs following the K-Radar (kradar) training protocol. During ASF training, we freeze all sensor-specific encoders to ensure that only the ASF components and detection head are learned. Specifically, ‘UCP’ and ‘CASAP’ are the ASF components updated during ASF training, which denote unified canonical projection and cross-attention across sensors along patches, respectively. Details for each encoder are as follows:
@@
Camera. Camera-based sensor encoders for autonomous driving typically require transforming front-view (FV) images into a bird’s-eye-view (BEV) format for 3D object detection. Current approaches can be broadly categorized into two methods: (1) learnable BEV transformations such as Lift-Splat-Shoot (LSS) (lss, bevfusion_icra, bevfusion_nips), or (2) maintaining FV while using transformer decoders (petr) that incorporate depth information as positional embeddings in BEV space. Although methods like PETR (petr) have improved 3D object detection by injecting depth embeddings, they inevitably increase computational complexity due to the additional depth dimension. These two approaches form the foundation of camera-based fusion (lxl, rcfusion, bevfusion_icra, bevfusion_nips, cmt), commonly seen in deeply coupled fusion (DCF) and sensor-wise cross-attention fusion (SCF).
@@
To develop a computationally efficient method, we opt for the LSS approach and adopt BEVDepth (bevdepth), which supervises the LSS depth estimation using LiDAR point clouds. For encoding images, we use the Swin-S model of SwinTransformer as the 2D backbone, achieving 83% top-1 accuracy on ImageNet (imagenet). We resize front camera images to $704\!\times\!256$ pixels for encoding. The encoded 2D FMs are projected into BEV FMs via LSS with a grid size of 0.2 m. To match the 0.4 m grid size used in K-Radar, we add a convolutional layer with stride 2. During training, we employ the depth information projected from LiDAR point clouds onto the image and, inspired by (mono_distil), we use the LiDAR backbone’s (SECOND) BEV FMs for distillation. As shown in Fig. 3, camera images have limited utility in adverse weather, so we train the camera encoder only on normal or overcast scenes (sequences 1–20) from K-Radar.
@@
A limitation of this approach is that promising stereo or multi-view detection models cannot be fully leveraged, because the K-Radar dataset (kradar)—the only one offering camera, LiDAR, and 4D Radar data in adverse weather—provides object annotations only for the front view. Moreover, unlike the KITTI dataset with its longer stereo baseline of 0.54 m, K-Radar uses waterproof ZED cameras with a baseline of just 0.12 m, limiting stereo-based detection for distant objects. Consequently, we employ a monocular camera method despite inherent performance constraints.
@@
LiDAR. LiDAR provides precise depth information in the form of dense point clouds, often used as a reference for labeling other sensors (kradar, auto_label). We adopt the SECOND (second) model as the LiDAR encoder, which is open-source in K-Radar and widely used in various benchmarks. Leveraging LiDAR’s high spatial resolution, we set the voxel size to 0.05 m and apply multiple layers of sparse convolution to generate BEV FMs with a 0.4 m grid size.
@@
4D Radar. Unlike LiDAR, which uses Time-of-Flight (ToF) measurement, Radar detects spatial information through signal processing, which can introduce sidelobe noise. Traditional 3D Radar typically filters out noisy measurements to produce a sparse point cloud focusing on object presence (radar_tutorial). In contrast, 4D Radar captures additional elevation information, yielding higher-density point clouds (dual). This property is exemplified by K-Radar (kradar, enhanced_kradar). We adopt RTNH (kradar), the baseline provided in K-Radar, as our 4D Radar encoder, following the original network settings and design.
@@ A.2 Detection Head
Following the K-Radar baseline, we use an SSD detection head (ssd, complexyolo) for the 3D object detection task. Our implementation uses two anchor boxes per object class, oriented at $0^{\circ}$ and $90^{\circ}$. Each anchor box is parametrized by eight values: the 3D center coordinates ($x_{c}$, $y_{c}$, $z_{c}$), 3D size ($x_{l}$, $y_{l}$, $z_{l}$), and orientation ($\cos(\text{yaw})$, $\sin(\text{yaw})$), as in (complexyolo).
To address the severe class imbalance in object detection (negative samples far outnumber positives), we use the focal loss (focal) for classification. For bounding box regression, we use smooth L1 loss to measure differences between predictions and ground truth, improving training stability. During inference, we select the class with the highest logit score for each proposal, then apply non-maximum suppression to remove redundant overlapping detections, producing the final predictions.
@@ B.1 Additional t-SNE Visualizations
Figure 5: Visualization of feature representations through t-SNE (tsne) at different stages of the proposed ASF for the ‘Bus or Truck’ class. ‘UCP’ and ‘CASAP’ denote unified canonical projection and cross-attention across sensors along patches, which are the two main components of the proposed ASF.
This section provides additional feature visualizations to complement those in Fig. 2, which only shows results for the ‘Sedan’ class. Fig. 5 presents t-SNE (tsne) visualizations of the ‘Bus or Truck’ class at different stages of the ASF. These visualizations further demonstrate how ASF unifies feature representations from different sensors.
Consistent with Fig. 2, we observe three main trends in the t-SNE visualizations. First, the initial features from the sensor-specific encoders are scattered without clear alignment across different sensors. Next, after UCP, sensor-specific features become better aligned with the fused features. Finally, following CASAP, features from available sensors form cohesive clusters (in this case, LiDAR and 4D Radar).
These findings underscore the effectiveness of ASF in creating a unified canonical space for multi-sensor features. Note that in K-Radar, the ‘Sedan’ class has 37,213 instances, whereas the ‘Bus or Truck’ class has 7,449 instances, leading to sparser data in Fig. 5 compared to Fig. 2.
@@ B.2 Details of t-SNE Visualizations
To visualize object-specific features in each sensor’s feature map (FM), we use RoIAlign (roialign) to crop features within bounding boxes. RoIAlign extracts fixed-size FMs, making it suitable for t-SNE (which requires consistent dimensions). We then visualize these cropped features in the unified canonical space. For those extracted before projection (where channel dimensions may differ), we downsample as needed to ensure consistent feature dimensions.
Appendix C Analysis of Attention Scores across Sensors
Table 5: Attention scores ratios [%] across different weather conditions
Sensor Normal Overcast Fog Rain Sleet Light Snow Heavy Snow
Camera 11.6 11.1 11.3 11.1 11.1 10.9 11.0
LiDAR 39.5 37.5 37.5 37.3 35.9 36.3 34.7
4D Radar 48.9 51.4 51.2 51.6 53.0 52.8 54.3
Table 6: Attention scores ratios [%] across different distance ranges
Sensor 0-8m 8-16m 16-24m 24-32m 32-40m 40-48m 48-56m 56-64m 64-72m
Camera 9.8 10.1 10.4 10.9 11.1 11.3 11.9 12.7 13.4
LiDAR 48.4 40.2 37.2 36.0 34.9 34.8 35.0 35.5 36.0
4D Radar 41.8 49.7 52.4 53.1 54.0 53.9 53.1 51.8 50.6
Tables 5 and 6 show the attention score ratios for each sensor according to weather conditions and distance from the ego-vehicle. Intuitively, a higher attention score ratio indicates greater utilization of the corresponding sensor in ASF. Note that the sensor attention maps (SAMs) in Fig. 3 visualize the attention scores for each sensor.
As shown in Tab. 5, the attention score ratios of camera and LiDAR decrease under severe adverse weather conditions such as sleet and heavy snow (camera: 11.6% to 11.0%, LiDAR: 39.5% to 34.7%), while 4D Radar shows the opposite trend, increasing from 48.9% to 54.3%. This demonstrates that 4D Radar is more robust than camera and LiDAR in adverse weather conditions and provides available measurements when other sensors degrade.
Furthermore, as shown in Tab. 6, LiDAR’s attention score ratio decreases as distance increases (from 48.4% at 0-8m to 36.0% at 64-72m), while camera’s attention score increases (from 9.8% to 13.4%). This indicates that ASF relies more on camera measurements at longer ranges where LiDAR becomes sparse. In contrast, 4D Radar maintains a relatively consistent attention ratio (41.8% to 50.6%) regardless of distance. Notably, 4D Radar shows the highest absolute attention scores compared to camera and LiDAR across most conditions. While camera lacks depth information and LiDAR provides very sparse measurements in 3D space, 4D Radar leverages dense tensor data, which enables more reliable feature extraction for sensor fusion.
Appendix D Additional Detection Results
Figure 6: Qualitative results of ASF with C+L+R (C: Camera, L*: damaged LiDAR, R: 4D Radar). We show results under (a–b) normal and (c–d) adverse weather conditions. Each grid shows the camera image, LiDAR point cloud, 4D Radar tensor, and a sensor attention map (SAM) illustrating attention scores from CASAP (except for (a) and (c), which only show sensor data for ground truth visualization). In the SAM, red, green, and blue represent attention to camera, LiDAR, and 4D Radar, respectively. For example, a predominantly blue SAM indicates that 4D Radar has the highest attention in that scene. The bottom-left corner of each grid shows attention score proportions (C/L/R[%]). Yellow and red boxes denote predictions and ground truth, respectively.
This section presents additional detection results for ASF. First, we provide qualitative results demonstrating ASF’s robustness when the LiDAR sensor is damaged, shown in the same road environment as Fig. 3, along with additional results from diverse road scenarios. Second, we demonstrate the robustness of ASF (C+L+R) with respect to random seeds by presenting experimental results on the K-Radar (kradar) benchmark v1.0 using 10 different random seeds (0 to 8 and 2025), complementing Tab. 1 which reports results only for seed 2025. Third, we evaluate alternative camera backbone architectures, including stereo camera-based methods, to assess the impact of improved camera encoders on the overall fusion performance. Finally, we provide comprehensive quantitative performance on K-Radar benchmark v2.0, including $AP_{3D}$ and $AP_{BEV}$ for both ‘Sedan’ and ‘Bus or Truck’ classes at IoU=0.3 and 0.5, complementing Tab. 3 which reports only $AP_{3D}$ at IoU=0.3.
Figure 7: Qualitative results of ASF for various sensor combinations. We show results low-light nighttime roads. Each grid shows the camera image, LiDAR point cloud, 4D Radar tensor, and a SAM illustrating attention scores from CASAP. In the SAM, red, green, and blue represent attention to camera, LiDAR, and 4D Radar, respectively. The bottom-left corner of each grid shows attention score proportions (C/L/R[%]). Yellow and red boxes denote predictions and ground truth, respectively, which are displayed in all panels. Dotted triangles indicate miss detections and are marked in the SAM. Note that predictions are visualized on all sensor data, even when a sensor is not employed for detection (e.g., predictions from L+R are also visualized on the camera image).
Figure 8: Qualitative results of ASF for various sensor combinations. We show results under adverse weather conditions. Each grid shows the camera image, LiDAR point cloud, 4D Radar tensor, and a SAM illustrating attention scores from CASAP. In the SAM, red, green, and blue represent attention to camera, LiDAR, and 4D Radar, respectively. The bottom-left corner of each grid shows attention score proportions (C/L/R[%]). Yellow and red boxes denote predictions and ground truth, respectively, which are displayed in all panels. Dotted triangles and circles indicate miss detections and false alarms, respectively, and are marked in the SAM. Note that predictions are visualized on all sensor data, even when a sensor is not employed for detection (e.g., predictions from L+R are also visualized on the camera image).
D.1 Additional Qualitative Results
@@
Damaged LiDAR. When LiDAR surfaces are damaged or contaminated, LiDAR measurements can be lost (lidar_contamination). In adverse conditions such as heavy snow, accumulation on the LiDAR sensor can block forward measurements, as illustrated by the red dashed line in Fig. 6-(d), even though vehicles are present. We also simulate total loss of forward LiDAR returns from surface damage (e.g., broken sensor), as seen in Fig. 6-(b). Despite missing LiDAR data, ASF still detects all objects using only camera and 4D Radar inputs. As reported in Tab. 3, ASF with C+L+R often performs similarly or even better than C+R alone. Compared to Fig. 3-(b), the sensor attention map (SAM) in Fig. 6-(b) confirms the increased reliance on camera or 4D Radar when LiDAR is unavailable.
Additional Results Visualization. Fig. 7 shows the inference results for multiple vehicles in low-light road environments. When detection is performed using camera-only in low-illumination conditions, numerous miss detections occur, as shown in Fig. 7-(f). Additionally, as shown in Fig. 7-(h), in the case of 4D Radar, one of the distant vehicles aligned side by side is covered by the reflection of another vehicle, resulting in miss detection. Furthermore, the leftmost vehicle experiences miss detection as the radio waves are blocked by the road barrier. In the case of Fig. 7-(a), which utilizes all of 4D Radar, LiDAR, and camera, stable and accurate detection is possible without miss detections.
Fig. 8 shows the inference results for multiple vehicles in adverse weather conditions. As shown in Fig. 8-(e), when only LiDAR is utilized, nearby vehicles can be detected due to the presence of measurements, but distant vehicles experience miss detection due to the absence of measurements. In contrast, as shown in Fig. 8-(h), 4D Radar has measurements regardless of distance, enabling stable detection without miss detections. Similar to Fig. 7-(a), in the case of Fig. 8-(a), which utilizes all of 4D Radar, LiDAR, and camera, stable and accurate detection is possible without miss detections and false alarms.
D.2 Additional Quantitative Results
Table 7: Additional performance comparison of ASF (C+L+R) under various weather conditions on the K-Radar (kradar) benchmark v1.0. We report both $AP_{BEV}$ and $AP_{3D}$ at IoU = 0.3 for the ‘Sedan’ class, showing mean and standard deviation across 10 random seeds (0 to 8 and 2025).
Metric Total Normal Overcast Fog Rain Sleet Light snow Heavy snow
$AP_{BEV}$ 88.64 88.24 90.29 98.97 88.93 86.24 89.25 78.44
$\pm$0.12 $\pm$0.12 $\pm$0.09 $\pm$0.27 $\pm$0.17 $\pm$2.85 $\pm$0.11 $\pm$0.16
$AP_{3D}$ 87.48 86.96 89.94 90.67 88.28 80.46 88.92 76.03
$\pm$0.21 $\pm$0.17 $\pm$0.20 $\pm$0.08 $\pm$0.20 $\pm$0.21 $\pm$0.22 $\pm$2.33
Multiple Random Seeds. Tab. 7 presents the mean and standard deviation of ASF (C+L+R) performance across 10 random seeds on the K-Radar benchmark v1.0, demonstrating the robustness and generality of our approach. The low standard deviations (mostly below 0.3) indicate stable performance across different random initializations. Notably, the sleet and heavy snow condition shows higher variance ($\pm$2.85 for $AP_{BEV}$ and $\pm$2.33 for $AP_{3D}$, respectively), suggesting this weather condition presents more challenging and variable detection scenarios. The consistently high performance in Fog conditions (98.97 $AP_{BEV}$ and 90.67 $AP_{3D}$) with minimal variance demonstrates ASF’s particular effectiveness in low-visibility scenarios where 4D Radar excels.
Table 8: Performance comparison of monocular and stereo camera-based networks on the K-Radar (kradar) benchmark v2.0. We report both $AP_{BEV}$ and $AP_{3D}$ at IoU = 0.3 for the ‘Sedan’ and ‘Bus or Truck’ classes. We note that camera performance is evaluated on sequences 1 to 20, which consist of normal weather conditions.
Metric Class BEVDepth (Mono) DSGN++ (Stereo) ASF (BEVDepth) ASF (Stereo)
$AP_{BEV}$ Sedan 19.9 24.3 82.5 82.4
Bus or Truck 13.4 24.8 70.2 67.2
$AP_{3D}$ Sedan 17.8 20.8 79.4 79.3
Bus or Truck 11.5 22.5 60.4 59.5
Stereo Camera. As shown in Tab. 8, the stereo camera-based DSGN++ (dsgnp) achieves improved performance on K-Radar (24.8% and 22.5% $AP_{3D}$ for Sedan and Bus or Truck, respectively) compared to the monocular camera-based BEVDepth (bevdepth). We integrate DSGN++ as ASF’s camera-specific encoder. However, as shown in Tab. 8, ASF using DSGN++ shows comparable performance to ASF using BEVDepth. This may be attributed to two factors: (1) the detection performance of LiDAR and 4D Radar is significantly superior to that of the camera, and (2) K-Radar’s ZED 2i camera has a 12cm baseline compared to KITTI’s 54cm, limiting the improvement potential of stereo-based methods.
Performance on Benchmark v2.0. Tab. 9, 10, 11, and 12 present comprehensive evaluations of ASF performance across various sensor configurations and weather conditions on the K-Radar benchmark v2.0. Results demonstrate that multi-sensor fusion approaches consistently outperform single-sensor methods. The C+L+R configuration achieves the highest performance in most scenarios for both ‘Sedan’ and ‘Bus or Truck’ classes, with C*+L+R often ranking second, which indicates that a damaged camera does not significantly affect performance.
Performance analysis across weather conditions reveals that 4D Radar-inclusive combinations (C+L+R, L+R) exhibit superior performance in adverse weather (demonstrating 4D Radar’s effectiveness under challenging environmental conditions), particularly in fog and snow. While camera-LiDAR fusion (C+L) occasionally outperforms in normal and overcast conditions, its performance deteriorates significantly in precipitation.
Table 9: Performance comparison of ASF under various sensor combinations evaluated on the K-Radar (kradar) benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{3D}$ at IoU = 0.3 under various weather conditions for the ‘Sedan’ and ‘Bus or Truck’ classes. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained with R+L+C.